-14/9 / -2 5/6
and
3/2 / 1 5/8
thank you soo much!

Answer by stanbon(75887)

(-14/9) / (-2 5/6)
------
= (-14/9) / [-(6*2+5)/6]
---
= (-14/9) / [-17/6]
---
Invert and multiply:
= (-14/9)*(-6/17)
---
Reduce where you can to get:
= (-14/3)*(-2/17)
---
= 28/51
===========================
(3/2) / (1 5/8)
----
= (3/2) / (13/8)
---
= (3/2)*(8/13)
---
= 12/13
==============
Cheers,
Stan H.